A hollow spherical shell has inner diameter 2 cm and outer diameter 4 cm.Determine the volume of the shell.
Answer by SATCHI76(2)

Here R= 2 cm, r= 1 cm
Volume =(4/3). Pie. (R^3-r^3)
=(4/3). Pie. (2^3-1^3)
=(4/3). Pie. (8-1)
=(4/3). Pie. 7
=(4/3). 22
= 88/3 cm^3